OpenStudy (anonymous):

find the fourierbseries representation f(x)=x+1 for -1

OpenStudy (anonymous):

Do you know the Fourier series equations? What is a Fourier series?

OpenStudy (anonymous):

\[a_k = \frac{1}{\pi}\int\limits_{-\pi}^{\pi}f(x)\cos(kx) dx\]\[b_k = \frac{1}{\pi}\int\limits_{-\pi}^{\pi}f(x)\sin(kx) dx\]\[f(x) = \frac{a_0}{2}+ \sum _{n=1}^{\infty}a_k \cos(kx)+b_ksin(kx)\]

OpenStudy (anonymous):

\[a_0 =\frac{1}{\pi}\int\limits_{-\pi}^{\pi}f(x)\cos(0x) dx=\frac{1}{\pi}\int\limits_{-\pi}^{\pi}f(x) dx\]

OpenStudy (anonymous):

the above formulas work for a function f(x) with period 2pi. Thus it is required for the function to be defined in the interval (-pi, pi). In your question, nothing has been said about the period. Also, the given function is defined only for " 1<equal x<equal 1". so you'' have to modify the limits of integration and take only those values which are allowed by the domain.

OpenStudy (anonymous):

ok but can you please complete it for me.... i tried solving but i got many errors please help

OpenStudy (anonymous):

@akitav , your equations are incorrect. @GIL.ojei please refer to http://www.sosmath.com/fourier/fourier1/fourier1.html , and http://www.sosmath.com/fourier/fourier2/fourier2.html for the equations with limits [-L,L] and examples using those limits.

OpenStudy (anonymous):

ibut it is the right question f(x)=(x+1) -1\[\le x \le1\]

OpenStudy (anonymous):

You need to calculate your a0, an, and bn constants first. Show me what you have so far.

OpenStudy (anonymous):

i got 1/pi -2sinpix/pi +1/sinpix/pi ..... please solve and show me what you got if possible, please find \[\int\limits_{-1}^{1}(x+1)sinx \] for bn because an is =0.. show work

OpenStudy (anonymous):

That's not right, let's solve a0 first because it's easy. Show me your steps. The equation is \[a _{0}=\frac{ 1 }{ 2L }\int\limits_{-L}^{L} f(x)dx\]

OpenStudy (anonymous):

ok. please solve and let me watch and learn.. please it is an exam question and do not want to do any mistake ,,, please

OpenStudy (amorfide):

stating it is an exam question we can't help you, also it was explained very well